THE FOLLOWING INFORMATION IS USED FOR QUESTIONS 8 – 10.

Quiz Company provided the following information for the preparation of the cash budget for June.

The cash balance of June 1 is expected to be $3,763.

Projected sales are as follows.

April

May

June

Cash sales

$10,000

$18,000

$18,600

Credit sales

28,900

35,000

54,000

Credit sales are collected over a three-month period: 40% in the month of sale, 30% in the first month after the sale, 20% in the second month after the sale. Sales collected in the second month after the sale second month after the sale are subject to a 2% late fee that customers pay in addition to the amount owed. The remaining sales are uncollectible.

Inventory purchases average 64% of a month’s total sales. Purchases are all on credit and are paid 20% in the month of purchase and 80% in the month following the purchase.

Salaries and wages are $11,750 per month and are all paid during the month.

Other cash expenditures during the month are expected to be $6,780.

(Round all calculations to the nearest dollar.)

8. Determine cash collections for June.

Note: Give your answer using dollar signs and commas but no decimal points (cents).

Example: $12,345

9. Determine the total cash disbursements for June.

10. Determine the ending cash balance for June.

1. Quiz Company reported the following information for the last year.

Sales

$4,000,000

Cost of goods sold

     2,100,000

Gross margin

1,900,000

Operating expenses

    1,100,000

Operating income

800,000

Income taxes (40%)

       320,000

Net income

$480,000

Quiz Company began the year with operating assets of $2,700,000. At year end, operating assets were $2,300,000. Quiz Company’s required rate of return was 15%, and the cost of capital was 12%. Total capital employed during the year was $1,400,000.

Determine residual income (RI) for the year.

Note: Give your answer using dollar signs and commas but not decimal points (cents).

Example: $12,345

2. Determine economic value added (EVA) for the year.
Answer #1
(8) Determine the Cash collection for June
$
Cash Sales for the June            18,600
Credit Sales received in june for
June - (54000 x 40%)            21,600
May - (35000 x 30%)            10,500
April - (28900 x 20%) + (28900 x 30% x 2%)              5,953
Cosh collection in June Month            56,653
(9) Determine the total Cash disbursement for June
$
Total Sales for June (18600 + 54000)            72,600
Purchase inventory (72600 x 64%)            46,464
Paid 20% in same month (46464 x 20%) (a)              9,293
Paid 80% of May (33920 x 80%) (b)            27,136
Cash paid for inventory purchase (a+b)            36,429
Salaries & Wages            11,750
Other expenses              6,780
Total cash disbursement            54,959
* May Purchase is (18000+35000)x64% =            33,920
(10) Cash Budget for June
$
Beginning Cash Balance              3,763
Cosh collection            56,653
Cash disbursement          -54,959
Ending Cash Balance              5,458
(1) Residual Income = ( Operating Income - (Operating Assets* Required Rate of return)
      Residual Income = ( 800000 - (2700000* 15%) =         $        3,95,000
2) EVA = Net operating profit after tax - (Total Assets - Total Liability) x WACC
     EVA = 480000 -(1400000*12%) =          $        3,12,000